A triangle has a perimeter of 84 inches. its shortest side measures 9 inches shorter than its middle side,and its longest side measures 6inches longer than its middle side. find the lengths of the triangle three sides
Answer by dewasmaharaja(5)

Let three sides of the triangle be
l for long
m for middle
s for short
Then,
l + m + s = 84 ...(a)
s = m - 9 ...(b)
l = m + 6 ...(c)
Substituting equation (b) and (c) in (a),
m - 9 + m + m + 6 = 84
3m = 87
Therefore, m = 87/3 = 29
From (b), s = 29 - 9 = 20
From (c), l = 29 + 6 = 35